OpenStudy (anonymous):

Find the slope and y-intercept of the line. 14x + 4y = 24

OpenStudy (anonymous):

Subtract both sides with 14

OpenStudy (uri):

Formula :y=mx+b

OpenStudy (anonymous):

Exactly so you gotta take the number from the x and subtract both sides

OpenStudy (anonymous):

meh puedas dar la repuesta ?

OpenStudy (anonymous):

Lol the slope is the number with the x it's always x and the y intercept is the number next too y

OpenStudy (anonymous):

Slope means x and y intercept means y, you get what I sayin?

OpenStudy (anonymous):

dame la repuesta amol :) real tawlk idont get math

OpenStudy (anonymous):

X+-10=24

OpenStudy (anonymous):

You in 9th cuz this algebra I gotta take ore calculus now, I'm tight

OpenStudy (anonymous):

im actually in 12th lol if i finish my online work i graduate on aug. 5th i juss dont get math

OpenStudy (anonymous):

and is the answer 14;24 ?

OpenStudy (anonymous):

Oh dame, I'm in 11 now I'm doin summer work so I coiled be out this joint and live next too kendo kaponi!! Lol nah I'm jokin but i gotta house near gringo if you know who that is

OpenStudy (uri):

Subtract 14x from both sides. 4y = 24 - 14x Divide by 4 y = 6 - 7/2 x

OpenStudy (uri):

Shift the 6 on right side. (-7/2)x + 6

OpenStudy (anonymous):

Yo ima stop helping you with math, cuz you guns probably get that sh it wrong

OpenStudy (uri):

-7/2x is the slope and 6 is y intercept.

OpenStudy (anonymous):

dique gringo lol if it was next to arcangel id be like woow! LOL and thanks uri!

OpenStudy (anonymous):

and why you say that?

OpenStudy (anonymous):

Lol I didn't say it to impress I was serious my aunt is right across from him, and I'm movin with her when done with high school

OpenStudy (anonymous):

And arcangel ain't all that he ugly as F$$$$ I met him and his wife,

OpenStudy (anonymous):

Yo all dem n$$$ be in Hato Rey

OpenStudy (anonymous):

Uri if you divide 14 and 4 isn't it 3.5

OpenStudy (anonymous):

idc if he uglu his music is still too good ma dude

OpenStudy (anonymous):

Lol yea but ima girl and I thought he was hot untill you see him in person

OpenStudy (anonymous):

where you located at?

OpenStudy (anonymous):

You go first and answer that question

OpenStudy (anonymous):

i was born in rahway, Nj but now im living in NC

OpenStudy (anonymous):

I was born in NYC the Bronx and now I just moved to Florida

OpenStudy (uri):

No one comments after me ================ This is not a chat section I'm getting notifications..This is your last warning,Thanks.Go Pm eachother.

OpenStudy (anonymous):

I got Ny tated on me lol

OpenStudy (uri):

NO ONE COMMENTS AFTER ME DUDE.

OpenStudy (anonymous):

Lol dame you actually been there, I'm on vacation now ova here and it's wack